1/3

Дмитрий Николаевич Садовников,

Sten’ka Razin

tradukita de Tr. Jacob Robbins

1. From beyond the wooded island To the river wide and free Proudly sailed the arrow-breasted Ships of Cossack yeomanry.

2. On the first is Stenka Razin With a princess at his side, |: Drunken, holds a marriage revel With his beautiful young bride. :|

3. But behind them rose a whisper, “He has left his sword to woo; |: One short night, and Stenka Razin Has become a woman too!” :|

4. Stenka Razin hears the jeering Of his discontented band, |: And the lovely Persian princess He has circled with his hand. :|

5. His black brows have come together As the waves of anger rise, |: And the blood comes rushing swiftly To his piercing, jet-black eyes. :|

…

2/3

6. “I will give you all you ask for, Life and heart, and head and hand,” |: Echo rolls the pealing thunder Of his voice across the land. :|

7. “Volga, Volga, mother Volga, Deep and wide beneath the sun, |: You have never seen a present From the Cossack of the Don. :|

8. And that peace might rule as always All my free-born men and brave, |: Volga, Volga, mother Volga, Volga, make this girl a grave.” :|

9. With a sudden, mighty movement, Razin lifts the beauty high, |: And he casts her where the waters Of the Volga move and sigh. :|

10. Now a silence like the grave sinks To all those who stand to see, |: And the battle-hardened Cossacks Sink to weep on bended knee. :|

11. “Dance, you fool, and men, make merry! What has got into your eyes? |: Let us thunder out a chanty Of a place where beauty lies.” :|

…

3/3

12. From beyond the wooded island To the river wide and free, |: Proudly sail the arrow-breasted Ships of Cossack yeomanry. :|

Traduko de la Rusa poemo “Из-за острова на стрежень (Стенька Разин)” de Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883) en la Anglan de Tr. Jacob Robbins en 1921.

Arg-478-1039 (2010-03-09 17:06:16)

Tiu ĉi angligo troviĝas en http: // ingeb. org/ songs/ stenkara. html .